OpenStudy (anonymous):

Y=10 when x=5 find y when x=2 solve inversely

OpenStudy (anonymous):

y = 10 when x = 5 implies x and y share a relationship stated as x = 2y therefore for x = 2 y = 2 *2 = 4

OpenStudy (anonymous):

sorry relation is y = 2x

OpenStudy (wolf1728):

If y = 10 when x = 5 then x=y/2 yes it is cutvclrc